### v1

@@ 4 Photon surface analysis
Once we have established that the photon surface uniquely extends in the interior spacetime as a null hypersurface, let us study its qualitative behavior, to understand whether it covers the singularity developing in the LTB model.
@@
and defined in a left neighborhood of $x_{b}$, possibly extendable up to the regular centre $x=0$. That is possible because $r_{x}>0$ and therefore $x(t)$ can be inverted. In this way we can exploit all the techniques known in literature and based on comparison theorems in ODE [23, 22]. These techniques are based on a remarkable property [26] of the apparent horizon curve $t_{h}(s)$ to be a subsolution of the null geodesic equation $t^{\prime}(x)=r_{x}(t(x),x)$, i.e.
@@
Then, the photon surface will cover the singularity completely (i.e. including the center) if one of the two following instances happens:
@@
We will observe that the situation varies depending on whether the central singularity is naked or not. The initial scenario arises when, in (12), we have $n=1,2$ or $n=3$ and $a>{2(26+15\sqrt{3})}/{27}$. These cases specifically align with those where we can identify a super-solution for (18) of the form $t_{\xi}=1+\xi x^{n}$, selecting $\xi$ such that $t_{\xi}(x)<t_{h}(x)$ in a right neighborhood of $x=0$. The star boundary $x_{b}$ will be chosen in this neighborhood.
Assume that it is possible to select $t_{0}$ as defined in (14) such that $t_{\xi}(x_{b})<t_{0}<t_{h}(x_{b})$. In this case, the continuation of the solution to the Cauchy problem (18) extends to smaller values of $x$, reaching $x=0$, which results in $t(x)\to 1$. Consequently, the extension of the photon surface originates from the central singularity. Moreover, consider a time $t^{*}\in]t_{\xi}(x_{b}),t_{0}[$, the radial null geodesic with initial condition $t(x_{b})=t^{*}$ will be extended backwards up to $x=0$, again with $t(x)\to 1$, giving rise to a photon (actually, an infinite number of photons, given the arbitrariness of $t^{*}$) emanating from the central singularity but escaping the photon surface.
@@
we only have to check that $t_{\xi}(x_{b})<t_{0}$. Using (14) we obtain that this condition is satisfied in a right neighborhood of $x=0$ if and only if $n=1,2$ or $n=3$ and $a>(2/3)^{3/2}$. Therefore the mass profiles ensuring a central naked singularity also imply that the photon surface extends up to the central singularity, but it is insufficient to cover it.
The situation described here is represented in Figures 3 and 4, where the parameters are set in order to have a naked central singularity.
@@
The curve $t_{\xi}(x)$ is a supersolution of (18) and $x_{b}$ chosen such that $t_{0}\in]t_{\xi}(x_{b}),t_{h}(x_{b})[$. The solution of (18) can be extended back for $x<x_{b}$: it cannot cross neither the subsolution $t_{h}(x)$ from below nor the supersolution $t_{\xi}(x)$ from above, then it must extend up to the central singularity $(x=0,t=1)$.
Conversely, when $n>3$ or $n=3$ with a sufficiently small $a$, resulting in a covered central singularity, the photon surface will extend back to the regular center, see Figure 5.
@@
If it were to reach the central singularity, it would be a null radial geodesic escaping the central singularity, violating the conditions as the central singularity itself is trapped in this context.
We can summarize the above discussion in the following
@@
In the marginally bound LTB collapse, the photon surface extends back to the central singularity if and only if this is naked. In this case, there exist null radial geodesics $t_{g}(x)$ emanating from the central naked singularity such that $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$, where $t(x)$ is the solution of (18) associated with the photon surface extension established in Theorem 2.
@@ 5 Discussion and conclusions
Our results indicate that the only physically consistent extension of the photon surface $r=3M$ into the collapsing cloud is given by a surface generated by outgoing null radial geodesics. This extension satisfies the photon surface equation and adheres to the junction conditions at the boundary with the Schwarzschild exterior. Importantly, the nature of the central singularity—whether it is covered or naked—plays a crucial role in determining the global structure of the photon surface. When the singularity is covered, the photon surface terminates at the regular center before the singularity forms. Conversely, if the singularity is naked, the photon surface can extend back all the way to the central singularity itself. However, in such cases, the surface is not sufficient to prevent the escape of photons from the singularity, highlighting its limited covering capability.
These findings, relying on purely analytical studies, refine previous contributions – in particular [2] – and contribute to the ongoing discussion about the role of photon surfaces in dynamical spacetimes and their potential connection to cosmic censorship, showing that photon surfaces, though geometric in nature, are deeply intertwined with the causal and global structure of the underlying spacetime. Future research may explore whether similar conclusions hold in more general collapse scenarios, including non-marginally bound cases or models with non-zero pressure, as well as in the presence of rotation or other deviations from spherical symmetry.

### v2

@@ 4 Photon surface analysis
Once we have established that the photon surface uniquely extends to the interior spacetime as a null hypersurface, let us study its qualitative behavior to understand whether it covers the singularity developing in the LTB model.
@@
and defined in a left neighborhood of $x_{b}$, possibly extendable up to the regular centre $x=0$. This is possible because $r_{x}>0$ and therefore $x(t)$ can be inverted. Consequently, we can exploit all the techniques known in the literature and based on comparison theorems in ODE [Mena:2001dr, Giambo:2002tp]. These techniques are based on a remarkable property [Giambo:2001wi] that the apparent horizon curve $t_{h}(s)$ is a subsolution of the null geodesic equation $t^{\prime}(x)=r_{x}(t(x),x)$, i.e.
@@
Then, the photon surface will cover the singularity completely (i.e. including the center) if one of the two following instances occurs:
@@
We will observe that the situation varies depending on whether the central singularity is naked or not. The initial scenario arises when, in (12), we have $n=1,2$, or $n=3$ and $a>{2(26+15\sqrt{3})}/{27}$. These cases specifically align with those where we can identify a super-solution for (18) of the form $t_{\xi}=1+\xi x^{n}$, selecting $\xi$ such that $t_{\xi}(x)<t_{h}(x)$ in a right neighborhood of $x=0$. The star boundary $x_{b}$ will be chosen in this neighborhood.
Assume that it is possible to select $t_{0}$ as defined in (14) such that $t_{\xi}(x_{b})<t_{0}<t_{h}(x_{b})$. In this case, the continuation of the solution to the Cauchy problem (18) extends to smaller values of $x$, reaching $x=0$, which results in $t(x)\to 1$. Consequently, the extension of the photon surface originates from the central singularity. Moreover, consider a time $t^{*}\in]t_{\xi}(x_{b}),t_{0}[$, the radial null geodesic with initial condition $t(x_{b})=t^{*}$ will be extended backward up to $x=0$, again with $t(x)\to 1$, giving rise to a photon (actually, an infinite number of photons, given the arbitrariness of $t^{*}$) emanating from the central singularity but escaping the photon surface.
@@
we only have to check that $t_{\xi}(x_{b})<t_{0}$. Using (14) we obtain that this condition is satisfied in a right neighborhood of $x=0$ if and only if $n=1,2$ or $n=3$ and $a>(2/3)^{3/2}$. Therefore, the mass profiles ensuring a central naked singularity also imply that the photon surface extends up to the central singularity, but it is insufficient to cover it.
The situation described here is represented in Figures 3 and 4, where the parameters are set to produce a naked central singularity.
@@
The curve $t_{\xi}(x)$ is a supersolution of (18) and $x_{b}$ chosen such that $t_{0}\in]t_{\xi}(x_{b}),t_{h}(x_{b})[$. The solution of (18) can be extended back for $x<x_{b}$: it cannot cross the subsolution $t_{h}(x)$ from below nor the supersolution $t_{\xi}(x)$ from above and then it must extend up to the central singularity $(x=0,t=1)$.
Conversely, when $n>3$ or $n=3$ with $a$ sufficiently small, resulting in a covered central singularity, the photon surface will extend back to the regular center, see Figure 5.
@@
If it were to reach the central singularity, it would represent a null radial geodesic escaping the central singularity, violating the condition that the central singularity itself is trapped in this context.
We can summarize the above discussion as follows.
@@
In the marginally bound LTB collapse, the photon surface extends back to the central singularity if and only if it is naked. In this case, there exist null radial geodesics $t_{g}(x)$ emanating from the central naked singularity such that $t_{g}(x)<t(x)$, $\forall x\in]0,x_{b}]$, where $t(x)$ is the solution of (18) associated with the photon surface extension established in Theorem 2.
@@ 5 Discussion and conclusions
Our results indicate that the only physically consistent extension of the photon surface $r=3M$ into the collapsing cloud is given by a surface generated by outgoing null radial geodesics. This extension satisfies the photon surface equation and adheres to the junction conditions at the boundary with the Schwarzschild exterior. Importantly, the nature of the central singularity—whether it is covered or naked—plays a crucial role in determining the global structure of the photon surface. When the singularity is covered, the photon surface terminates at the regular center before the singularity forms. Conversely, if the singularity is naked, the photon surface can extend all the way back to the central singularity itself. However, in such cases, the surface is not sufficient to prevent the escape of photons from the singularity, highlighting its limited covering capability. Previous studies have established that naked singularities create different lensing effects compared to black holes [Virbhadra:2022iiy]. Additionally, research by Ortiz et al. [Ortiz:2015rma] demonstrates differences in the shadow evolution between naked singularities and black holes. While both objects eventually appear similar to distant observers, their formation dynamics differ significantly, also due to the infinitely redshifted photons that escape from the central naked singularity to the external region – a phenomenon documented in earlier work [Dwivedi:1998ts, Giambo:2006vr]. Our findings in the present paper support these observations by showing that shadow evolution also varies in terms of photon surface extension, suggesting that understanding photon surface geometry is crucial for distinguishing these compact objects, as the formation and features of the shadow are determined by the whole underlying spacetime structure.
These findings, relying on purely analytical studies, refine previous contributions – in particular [Cao:2019vlu] – and contribute to the ongoing discussion about the role of photon surfaces in dynamical spacetimes and their potential connection to cosmic censorship, showing that photon surfaces, though geometric in nature, are deeply intertwined with the causal and global structure of the underlying spacetime. Future research may explore whether similar conclusions hold in more general collapse scenarios, including non-marginally bound cases or models with non-zero pressure, as well as in the presence of rotation or other deviations from spherical symmetry.
Acknowledgement.
The authors thank LM Cao and Y Song for their insightful discussions.